OpenStudy (anonymous):

Prove -1x-1=1 using axioms explain your steps

OpenStudy (mertsj):

We can't prove that because if x = 2 for example, it is false.

OpenStudy (anonymous):

no its not

OpenStudy (mertsj):

-1(2)-1=-3 NOT 1

OpenStudy (anonymous):

i think what this means is \[-1\times -1=1\]

OpenStudy (anonymous):

its a multiply symbol not an x

OpenStudy (mertsj):

Well satellite, you just proved my point.

OpenStudy (anonymous):

how'd i guess

OpenStudy (mertsj):

??

OpenStudy (anonymous):

-1 (multiplied by) -1 = 1 PROVE

OpenStudy (anonymous):

mertsj u ever did axioms?

OpenStudy (mertsj):

I've done axioms before but Satellite is much smarter than I so you should ask him.

OpenStudy (anonymous):

ur smart too dont wary :)

OpenStudy (anonymous):

what's axiom anyways?

OpenStudy (anonymous):

the building blocks of math it is the 1st ever rules that created math

OpenStudy (anonymous):

the point is that -1 is the additive inverse of 1, so you know how it behaves with addition, but not will multiplication the link between multiplication and addition is the distributive property

OpenStudy (anonymous):

pretty much

OpenStudy (anonymous):

that is, we know that for every a there exists a -a such that a + (-a)=0

OpenStudy (anonymous):

but that is only 1 rule though

OpenStudy (anonymous):

so we can do better than proving \[-1\times -1=1\] we can prove that \[-1\times a=-a\]

OpenStudy (anonymous):

and we do it via the distributive property first we note that \[a\times 0=a\times (0+0)=a\times 0+a\times 0\] the first equality because \[0+0=0\] as 0 is the additive identity, and the second because of the distributive property this means \[a\times 0=a\times 0+a\times 0\] and this implies (by subtracting \[a\times 0\] form both sides that \[a\times 0=0\]

OpenStudy (mertsj):

So maybe you could do something simple: The reciprocal of -1 is -1 since 1/-1 = -1. The product of reciprocals is 1 by the definition of reciprocals. So (-1)(-1)=1 because they are reciprocals and the product of reciprocals is 1

OpenStudy (mertsj):

What grade are you in Hawk?

OpenStudy (anonymous):

now that we know \[a\times 0=0\] we know that \[0=a\times (1+(-1))=a\times 1+a\times (-1)\] implies that \[-1\times a=-a\] by subtracting a from both sides

OpenStudy (anonymous):

im in grade 11

OpenStudy (mertsj):

So you'd better pay attention to Satellite. That little thing I did is more like grade 8 or 9

OpenStudy (anonymous):

and now since \[a\times -1=-a\] we know that \[-1\times -1=1\] since -1 is the additive inverse of 1

OpenStudy (anonymous):

in general if you want to show something that involves properties of addition and multiplication, you need the distributive law

OpenStudy (anonymous):

dont say subtracting please cause that doesnt exist say additive inverse

OpenStudy (anonymous):

ill hav to digest wat u just wrote...

OpenStudy (anonymous):

wait you said the 0=0+0 ( i know this is obvious) but wat rule is this?

OpenStudy (anonymous):

additive identity?

OpenStudy (anonymous):

yes, 0 is the additive identity, so \[a+0=a\] and therefore \[0+0=0\]

OpenStudy (anonymous):

yes but u wrote ax0=ax(0+0)?

OpenStudy (anonymous):

isnt that assuming that 0+0=0 and not using any of the laws?

OpenStudy (anonymous):

yes this is true because \[0+0=0\] so \[a\times 0=a\times (0+0)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\checkmark\]